(2 n+2 -2(2 n ))/2 2n+2 - simplify this plz this is diffrent

(2n+2-2(2n))/22n+2  - simplify this  plz this is diffrent

Grade:NULL

2 Answers

Arun
5 years ago
Dear Jatin
 
2n * 2² - 2 * 2n /2² * 2²n
= 2n( 4 -2) / 2²n *4
= 2/2n *4
= 1/2n * 2
= 1/2n+1

Regards
Arun (askIITians forum expert)
Jatin Soni
5 years ago
Can u do it on whitebord and send me plz. It is not clear like this. i am unable to get u. plz if u can